OpenStudy (cutiecomittee123):

PRE CALC HELP!!! Compute; (-3+2i)/(4+11i)

OpenStudy (unklerhaukus):

\[\frac{-3+2i}{4+11i}\\ =\frac{-3+2i}{4+11i}\times\frac{4-11i}{4-11i}\\ =\]

OpenStudy (cutiecomittee123):

Okay so basically we are doing cross multiplication here?

OpenStudy (unklerhaukus):

we are multiplying the denominator by its complex conjugate, this will realize the denominator

OpenStudy (cutiecomittee123):

alright makes sense

OpenStudy (anonymous):

i'm gonna go smack every math teacher that ever said "cross multiply"

OpenStudy (cutiecomittee123):

It helps me to remember when dealing with fractions

OpenStudy (unklerhaukus):

what does the new denominator become \[(4+11i)\times(4-11i)=?\]

OpenStudy (cutiecomittee123):

8+122i

OpenStudy (cutiecomittee123):

becasue we get 1-44i+44i-122i

OpenStudy (unklerhaukus):

try that again

OpenStudy (unklerhaukus):

\[(a+b)(a-b) = a^2-b^2\]

OpenStudy (anonymous):

and also (which is what you need ) \[(a+bi)(a-bi)=a^2+b^2\]

OpenStudy (unklerhaukus):

What happened to your SmartScore @satellite73?

OpenStudy (cutiecomittee123):

Im very lost now

OpenStudy (cutiecomittee123):

Like what numbers are a and b in this case?

OpenStudy (unklerhaukus):

a = 4 b = 11i

OpenStudy (cutiecomittee123):

(4+11i)(4-11i)=4^2+11i^2

OpenStudy (unklerhaukus):

almost \[(4+11i)(4-11i)=4^2-11i^2\] now use \(i^2=-1\)

OpenStudy (cutiecomittee123):

16+11=27*

OpenStudy (unklerhaukus):

whoops i mean \[(4+11i)(4-11i)=4^2-(11i)^2\]

OpenStudy (cutiecomittee123):

so its 16-11?

OpenStudy (unklerhaukus):

so \[4^2-11^2i^2=4^2+11^2\]

OpenStudy (cutiecomittee123):

Oh so then 16+122

OpenStudy (unklerhaukus):

yes, the denominator becomes 16 + 121

OpenStudy (cutiecomittee123):

137

OpenStudy (unklerhaukus):

So, you have\[\frac{-3+2i}{4+11i}\\ =\frac{-3+2i}{4+11i}\times\frac{4-11i}{4-11i}\\ =\frac{(-3+2i)(4-11i)}{4^2+11^2}\\ =\frac{(-3+2i)(4-11i)}{137}\\ =\]

OpenStudy (unklerhaukus):

Now simplify the numerator

OpenStudy (unklerhaukus):

\[(a+b)(c+d) = ac+ad+bc+bd\]

OpenStudy (unklerhaukus):

now a = -3, b = 2i, c = 4, d = -11i

OpenStudy (cutiecomittee123):

(-3(4)+-3(-11i)+2i(4)+2i(-11))/137

OpenStudy (unklerhaukus):

yeap, now simplify it

OpenStudy (cutiecomittee123):

(-12+33i+8i-22i)/137

OpenStudy (unklerhaukus):

good, keep going

OpenStudy (cutiecomittee123):

(-12+19i)/137

OpenStudy (unklerhaukus):

oh wait, the last term should have an i^2 (-3(4)+-3(-11i)+2i(4)+2i(-11i))/137 (-12+33i+8i-22i^2)/137

OpenStudy (unklerhaukus):

right?

OpenStudy (cutiecomittee123):

Yeah

OpenStudy (unklerhaukus):

and we know that i^2=−1 (-12+33i+8i+22)/137

OpenStudy (cutiecomittee123):

okay cool

OpenStudy (cutiecomittee123):

63i-12/137

OpenStudy (unklerhaukus):

not quite

OpenStudy (cutiecomittee123):

but 33+6+22=63 and the -12 goes with it

OpenStudy (cutiecomittee123):

-12+63i

OpenStudy (unklerhaukus):

you can only add the like terms, the 33i and 8i are like terms, and the -12 and +22 are like tems

OpenStudy (cutiecomittee123):

oh lol I forgot that the i^2 went away oops

OpenStudy (cutiecomittee123):

41i+10/137

OpenStudy (cutiecomittee123):

i think then you do 41i/137, 10/137

OpenStudy (unklerhaukus):

Either of those two forms the final result. Well done. \(\color{red}\checkmark\)

OpenStudy (cutiecomittee123):

sweet thanks for the help